OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

OpenStudy (anonymous):

Or if Yall can answer......... Simplify the given expression to rational exponent form and justify each step by identifying the properties of rational exponents used. All work must be shown. \[1/3\sqrt{x}^{-6}\]

OpenStudy (anonymous):

yes: remember that: \[a^m \times a ^m ... = a^{m +m...}\] \[x^3 \times x^3 \times x^3 \times x^3 = x ^{3+3+3+3} = x^{12}\] But he doesn't simplified it to the final/complete answer Ans2.) No his work wouln't be sameif he wrote like this: SInce,(simple laws of algebraic additions) \[x^3 + x ^3 + x^3 + x^3 => 4x^3\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

mention not bro :m +UP